# The Design Space of Lockfiles Across Package Managers

### v2

@@ 3.4.1 Lockfile Content
Package managers may include various information in lockfiles, incl. dependencies, checkums and additional details such as licenses. Figure 2 illustrates excerpts from lockfiles across the selected 7 package managers. Each of them contain an example representation of one dependency where the dependency versions, source links, checksums and indirect dependencies are highlighted in yellow, blue, green and pink respectively.
@@
Resolved package versions and checksums All of the studied package managers include resolved package versions in lockfiles. Checksums are also included in all lockfiles except in the Gradle lockfile. Gradle only records the groupId, artifactId, and version of the resolved dependencies which is a significant limitation.
Link to Source NPM and Cargo include a provenance link to source code. This is excellent for future rebuilding efforts [9]. In terms of naming, npm CLI uses the field resolved for source links, while Cargo uses source. In Go, the source can be inferred from the module name declaration itself, as we see from the link to the GitHub repository for the module text in 11. Gradle, pnpm, Poetry, and Pipenv do not include a direct link to the source of the downloaded packages.
@@
Additional metadata Some package managers include additional details such as language requirements (e.g. Poetry, npm), operating systems (e.g. npm, Cargo) or dependency scope (e.g. npm, Gradle). Moreover, npm CLI includes in the lockfile extra metadata such as license details and funding information, directly copied from package.json (5). Although recording as much information as possible for each dependency may seem a good choice, it also makes the lockfiles lengthier and harder to be reviewed by human developers.
Summary of lockfile content
All the package managers in this study record both direct and indirect dependencies in lockfiles, with all except Gradle also including dependency checksums. This information is essential for verifying the integrity of a build or for reproducing a build. However, among package managers, there are notable differences in the additional metadata included and the way this information is structured, which impact the compactness and readability of lockfiles.
@@
Figure 2: Example dependency representations extracted from lockfiles (or a mod file) across 7 package managers. Dependency versions are highlighted in yellow, source links in blue, checksums in green, and indirect dependencies in pink. Some details are omitted for brevity and clarity.
@@ 3.4.2 Lockfile Lifecycle
All 7 package managers in this study include a function to generate a lockfile, either by default or when explicitly enabled during dependency resolution. During subsequent project builds, the presence of a lockfile may change dependency resolution. At each of these steps of lockfile generation, dependency resolution, and enforcement, different package managers implement different policies, which we detail below. A summary of the key differences are presented in Table 4.
Lockfile generation. Package managers take different approaches to generate lockfiles. All package managers except for Gradle generate a lockfile by default. For Gradle, a lockfile is generated only when the –write-locks flag is enabled and lock states are explicitly defined in the dependency specification file.
Lockfile based dependency resolution. When a lockfile is present at the beginning of a build, it influences the dependency resolution process. Lockfile based dependency resolution is the property of respecting the locked states when resolving dependencies as long as they do not conflict with the dependency specification file. In npm, pnpm, Poetry, Gradle and Go, the build process respects the lockfile by default and only includes the locked versions. Conversely, with the build or install command, Cargo and Pipenv allow resolving new dependencies regardless of the locked versions, ignoring the lockfile. Afterwards, Cargo and Pipenv silently update the lockfile during the build. To make Cargo respect an exiting lockfile, developers have to use --locked or --frozen flags and for Pipenv they can use the --deploy flag or the pipenv sync command.
@@
Table 5: The most common types of information included in lockfiles
npm CLI $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (for each dependency), Dependency types, OS, engine, cpu, requirements, Licenses, Funding information
[…]
2613
[…]
Cargo $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (for each dependency), Language, OS, engine requirements
[…]
2613
[…]
2613
[…]
2613
Language requirements
[…]
2613
[…]
2613
[…]
2613
Dependency scopes
[…]
3.5 Results for RQ2: To what extent do open-source projects include lockfiles in version control?
@@
In total, we collect 4859 projects that include a dependency specification file from one of the 7 package managers we consider. Within the first six months of project creation, 3040 (62.6%) of these projects began committing lockfiles. For example, 43.4% have committed a package-lock.json, and 18% have committed a pnpm-lock.json within six months of the project creation date. Over time, 916(18.8%) more projects across all package managers have included lockfiles in version control. As of March 28, 2025, 81.4% of these projects have committed lockfiles to version control.
@@
For Go, most projects (92%) begin including go.sum within six months of project creation. The only exceptions are 4 projects: GoogleCloudPlatform/buildpacks, eolinker/apinto, go-admin-team/go-admin, and shalb/cluster.dev. These projects have 13, 48, 29, and 35 direct dependencies in their main modules, respectively. The developers of the project, go-admin-team/go-admin committed the lockfile at the early stages of the project but later decided to remove it from version control. The reason for removal is not documented in the public repository. The projects GoogleCloudPlatform/buildpacks and eolinker/apinto have go.sum files for some modules but do not contain one for the main module.
@@
Quantitative data alone does not fully reveal the exact reasons why developers choose to commit or omit lockfiles. To complement this first quantitative analysis across ecosystems, research questions 3 and 4 will be qualitative developer interviews.
@@ 4.2 Methodology for RQ3 and RQ4
To answer our 2 research questions on lockfile usage, we select a set of projects from each ecosystem, including some projects that publish their lockfile in the version control system and some projects that do not publish it. We then conduct semi-structured interviews with the top contributor of each project. In the following, we describe the methodology we follow to select the projects, contact developers, interview them and analyze their responses.
Participant selection: We aim to conduct interviews with developers who regularly contribute to open-source projects, have experience with managing dependencies, and maintain at least one project that uses one of the package managers under study.
@@
For each selected project, we gather publicly available contact information for the top contributor and send an interview invitation. If contact information is unavailable, or we do not receive a response within a week, we continue iterating through our randomly selected projects, until we reach our target of two or four participants per category, for a total of 14 responses. This approach allows us to select a diverse group of developers from various backgrounds and location, who have expertise in dependency management and who contribute to projects in various application domains.
Participant demographics: In total, we sent 59 invitations and received 14 responses: two npm/pnpm developers, two Cargo developers, two Poetry/Pipenv and two Go developers who commit lockfiles. We also received responses from two npm/pnpm developers, two Cargo developers, and two Poetry/Pipenv developers who do not commit lockfiles. We also interview one developer (P1) we were acquainted with through personal connections. In total we interview 15 developers and Table 7 presents the demographics of these participants. They work in 12 different domains and have an average of 8 years of experience. For anonymity, we denote each participant with the letter ”P” followed by an integer.
Interviews: We conduct 20 minute online interviews with 15 selected developers. We record the interviews upon consent from the participants. One interview (P8) was conducted in person at the participant’s request ; in this case, only written notes were taken. The participants were not provided any compensation for their participation.
@@
Interview response analysis: We use the Whisper speech recognition tool [18] to transcribe 15 interviews with available audio recordings. If audio recordings are not available, we refer to the written notes. Two authors carry out line-by-line coding for the transcribed interviews (or written notes). Consequently, the same two authors collaboratively create an affinity diagram, grouping codes into emerging themes inductively. The final affinity diagram consists of 20 sub-themes and 7 main themes related to general software development practices, lockfile benefits, developer expectations, and challenges.
@@
npm/pnpm P1 >{>}> 20 Yes 3k 110,892 320 Website development
@@ 4.3 Results for RQ3: What benefits do developers perceive in using lockfiles?
One of the main topics we discuss in the interviews with developers who commit lockfiles is their motivation for doing so. From their responses we identify the benefits they expect from using lockfiles. Further questions about debugging, code review, and security practices also reveal the direct and indirect benefits lockfiles offer. We identify benefits across four main areas: build determinism, integrity, transparency, and debugging.
@@ 4.3.1 Build determinism benefit
One of the core functionalities of lockfiles is supporting build determinism over time by locking dependency versions. A majority of developers who commit lockfiles (6/9) mention build determinism as the main reason for committing lockfiles. For instance, both Cargo developers who commit lockfiles state that they view the lockfile as a snapshot, and they are confident that the build will be identical every time they use the same lockfile (”Like I’m 100% certain that it’s going to be identical no matter when I go back to it.”, P6). The Poetry developer, P10, mentions the guarantee of build determinism that lockfiles provide across different machines (”the major benefit I see is that my state is exactly the same when I pull down the code on a new computer. So it’s just great for getting everything into the perfect, the same exact state that I was working with before.”)
Resolving dependency versions deterministically is particularly important in CI/CD pipelines to ensure reliable test results across environments and consistent builds before releases. P1 comments that they enforce and trust the lockfile in their CI/CD workflow (”When we go through the pipeline in order to test or to run our tests, typically we use npm ci to install the dependencies by trusting the package lock.”) P10 mentions that they always make sure every collaborator uses lockfiles in their release pipelines (”So like in your release pipeline, you want to make sure you’re using your lockfile, right? You don’t want to just be updating your whole dependency tree. You want to be using your lockfile.”)
@@ 4.3.2 Integrity benefit
All package managers in this study, except Gradle, record dependency checksums in their lockfile. These checksums allow the package manager to verify, during subsequent installations, that the recorded checksums match with the dependencies being downloaded from the registry. This verification helps detect potential package tamperings. P10 emphasizes its importance, stating, ”If you don’t have a hash of your dependency, you don’t know what your dependency is. It’s just that simple, right? You’re flying blind.
All developers trust that the checksums recorded in lockfiles are automatically validated by the package manager, and therefore, they do not carry out any additional checks. P1 note that (”We trust the fact that the checksums are being validated, but we’d never really investigate that.” Similarly, P10 explained ”My understanding is Poetry already does validation for the hashes. So when I go and I download, when I go and I sync my repository on a new machine, if the hash of the package doesn’t match what’s in my lockfile, it will throw me a warning.”. Poetry keeps a hash of the dependency specification file content in the lockfile and performs an additional hash validation when checking whether the dependency specification file has been updated and the lockfile is outdated. Poetry developers, such as P10, appreciate this additional check, (”So it will say your lockfile’s out of date. And then it will make me regenerate my lockfile, which I appreciate.”).
Identifying package tampering through integrity verification requires that releases remain immutable. In Go, when a package hosted on GitHub is modified after publication, such as through re-tagging or force-pushing, it can lead to checksum failures, as the version stored in the Go module proxy may differ from the modified release. A Go developer described this as a trade-off between the convenience of releasing and the security. (”it’s good on the one hand that you will always, like from a security perspective, make sure that you will always get the code that is basically stored on the proxy. And if you don’t have that, something is wrong and the checksum will give you an error. But then again, it can cause issues if you like want to re-release the same version. And yeah, it’s like convenience versus security”, P15)
@@ 4.3.3 Transparency benefit
Both Cargo developers who commit lockfiles mention that they prefer to verify pull requests that modify the dependency tree by re-generating the lockfile locally to ensure the changes are reproducible. (”Since I prefer to double check, I clone their fork locally, and I run cargo update, so that I’m sure that the lockfile is reflected on the actual changes, and, you know, maybe they had something malicious, so it’s better to check.”, P7.) P6 confirms it stating ”Typically what I will do is, I’ll pull their change into a branch. I will recreate the change that they did to the cargo.toml and then see if my lockfile changed in the exact same way.
Lockfiles, by exposing indirect dependencies, help developers monitor and manage dependency growth over time. P15 recall a time where they advised a collaborator to remove an unnecessary dependency ”So I said, okay, let’s not just import this whole dependency just for one Go file. Let’s just import the code over because it’s probably not changed anyways. And yeah, so usually I check the dependencies and basically question is this dependency needed or not”.

### v3

@@ 3.3.1 Lockfile Content
Package managers may include various information in lockfiles, dependencies, checksums, and additional details such as licenses. Figure 2 illustrates excerpts from lockfiles across the selected seven package managers. Each of them contains an example representation of one dependency where the dependency versions, source links, checksums, and indirect dependencies are highlighted in yellow, blue, green, and pink, respectively.
@@
Table 4: The most common types of information included in lockfiles
npm CLI
Indirect dependencies (for each dependency), Dependency types, OS, engine, CPU, requirements, Licenses, Funding information
pnpm
Indirect dependencies (under each dependency), OS, engine, CPU requirements
Cargo
Indirect dependencies (for each dependency), Language, OS, engine requirements
Poetry
Indirect dependencies (for each dependency), Language requirements, Descriptions
Pipenv
Language requirements
Gradle
Dependency scopes
Go
(in go.sum)
Resolved package versions and checksums All of the studied package managers include resolved package versions in lockfiles. Checksums are also included in all lockfiles except in the Gradle lockfile. Gradle only records the groupId (the namespace of the project), artifactId (the project name), and version (the project version) of the resolved dependencies, which is a significant limitation.
Link to source NPM and Cargo include a provenance link to the source of the downloaded package. Having a link to the source is useful for verifying that no tampering has occurred during the download from the package registry. In terms of naming, npm CLI uses the field resolved for source links, while Cargo uses source. In Go, the source can be inferred from the module name declaration itself, as we see from the link to the GitHub repository for the module text in 11. Gradle, pnpm, Poetry, and Pipenv do not include a direct link to the source of the downloaded packages.
@@
Additional metadata Some package managers include additional details such as language requirements (e.g. Poetry, npm), operating systems (e.g. npm, Cargo) or dependency scope (e.g. npm, Gradle). Moreover, npm CLI includes in the lockfile extra metadata such as license details and funding information, directly copied from package.json. One can note that not all of this metadata is included in the examples of Figure 2. This is because they show excerpts from real-world project lockfiles, where this information is not available in the dependency specification file, hence absent in the lockfiles as well. Although recording as much information as possible for each dependency may seem a good choice, it also makes the lockfiles lengthier and harder to be reviewed by human developers.
Summary of lockfile content All the package managers in this study record both direct and indirect dependencies in lockfiles. All except Gradle also include dependency checksums. However, among package managers, there are notable differences in the additional metadata included and the way this information is structured, which impact the compactness and readability of lockfiles.
@@
Figure 2: Example dependency representations extracted from lockfiles (or a mod file) across seven package managers. Each listing contains one extracted record corresponding to a direct dependency in a lockfile. Dependency versions are highlighted in yellow, source links in blue, checksums in green, and indirect dependencies in pink. Some details are omitted for brevity and clarity.
@@ 3.3.2 Lockfile Lifecycle
All seven package managers in this study include a function to generate a lockfile, either by default or when explicitly enabled during dependency resolution. At each of these steps of lockfile generation, dependency resolution, and enforcement, different package managers implement different policies, which we detail below. A summary of the key differences are presented in Table 5.
Lockfile generation. Package managers take different approaches to generate lockfiles. All package managers except for Gradle generate a lockfile by default. For Gradle, a lockfile is generated only when the --write-locks flag is enabled and lock states are explicitly defined in the dependency specification file.
Lockfile based dependency resolution. When a lockfile is present at the beginning of a build, it influences the dependency resolution process. Lockfile based dependency resolution is the property of respecting the locked states when resolving dependencies as long as they do not conflict with the dependency specification file. In npm, pnpm, Poetry, Gradle and Go, the build process respects the lockfile by default and only includes the locked versions. Conversely, with the default build or install command, Cargo and Pipenv resolves dependencies regardless of the locked versions, ignoring the lockfile. Afterwards, Cargo and Pipenv silently update the lockfile during the build. To make Cargo respect an existing lockfile, developers have to use --locked or --frozen flags and for Pipenv they can use the --deploy flag or the pipenv sync command.
@@
3.4 Methodology for RQ2
Next, we aim to analyze lockfile usage across package managers, by selecting a sample of public GitHub projects and collecting the lockfiles within them, if any.
Project selection: As the study subjects for RQ2, we collect popular public GitHub projects that use one of the seven package managers we consider in the study. When collecting GitHub projects, we aim to exclude personal toy projects and only target mature, widely-used projects [15]. To achieve this, we consider repositories created between September 30, 2019, and September 30, 2024, that have at least 42 stars, 10 contributors, and 300 commits. To ensure that we only select actively maintained repositories, we further exclude the projects that have not been updated within the last three months. For npm CLI and pnpm, we select repositories where the primary language is JavaScript or TypeScript. For Pipenv and Poetry, we include Python projects. For Cargo, Gradle, and Go, we target Rust, Java, and Go projects, respectively.
Lockfile identification: Similar to identifying the dependency specification file, we determine the presence of a lockfile by searching for the relevant lockfile files in the project directory. If a project has multiple lockfiles, we discard the project to prevent inconsistencies in the results and their interpretation. For example, in the JavaScript ecosystem, if a project contains both package-lock.json and pnpm-lock.json, we exclude it from the evaluation. We further disregard projects with lockfiles from package managers outside the selected seven package managers, such as yarn.lock from Yarn. For Go, we check for the existence of the go.sum file. It is our intuition that, for some developers, including a lockfile is related to the maturity of the project. As a project matures, the dependency tree may become more complex, which could lead developers to adopt a lockfile for better dependency management if they have not already done so. Therefore, we also collect the project creation date and the earliest lockfile commit date. Using this information, we derive the number of projects that began including lockfiles in version control within the first six months of project creation, as well as the total count of projects that currently have lockfiles.
3.5 Results for RQ2: To what extent do projects include lockfiles in version control?
@@
In total, we collect 4,859 projects that include a dependency specification file from one of the seven package managers we consider. Within the first six months of project creation, 3,040 (62.6%) of these projects began committing lockfiles. For example, 43.4% have committed a package-lock.json, and 18% have committed a pnpm-lock.json within six months of the project creation date. Over time, 916 (18.8%) more projects across all package managers have included lockfiles in version control. As of March 28, 2025, 81.4% of these projects have committed lockfiles to version control. This pattern suggests that developers are more likely to commit lockfiles as their projects mature.
@@
For Go, most projects (92%) begin including go.sum within six months of project creation. After six months only four projects do not commit lockfiles. Those four projects are: GoogleCloudPlatform/buildpacks, eolinker/apinto, go-admin-team/go-admin, and shalb/cluster.dev. These projects have 13, 48, 29, and 35 direct dependencies in their main modules, respectively. The developers of the project, go-admin-team/go-admin committed the lockfile at the early stages of the project but later decided to remove it from version control. The reason for removal is not documented in the public repository. The projects GoogleCloudPlatform/buildpacks and eolinker/apinto have go.sum files for some modules but do not contain one for the main module.
@@
These observations provide quantitative evidence confirming the findings of RQ1. First, it confirms that Gradle does not generate a lockfile by default and requires significant effort from developers to set it up, hence the absence of lockfiles. Second, Go requires minimal effort to generate and enforce lockfiles, hence the observed prevalence.
Quantitative data alone does not fully reveal the reasons why developers choose to commit or omit lockfiles. To complement this initial quantitative analysis across ecosystems, we conduct qualitative interviews with developers to better understand specific cases where lockfiles help them or do not fully satisfy their needs.
@@ 4.2 Methodology for RQ3 and RQ4
To answer our two research questions on lockfile usage, we select a set of projects from each ecosystem, including some projects that publish their lockfile in the version control system and some projects that do not publish it. We then conduct semi-structured interviews with the top contributor of each project. In the following, we describe the methodology we follow to select the projects, contact developers, interview them, and analyze their responses.
Participant selection: We aim to conduct interviews with developers who regularly contribute to open source projects, have experience with managing dependencies, and maintain at least one project that uses one of the package managers under study.
@@
For each selected project, we search the public GitHub profile of the top contributor and send an interview invitation if their email address is publicly available. As the main contributor of a mature project, we expect these developers to be knowledgeable about lockfiles and be familiar with dependency management best practices. When developers respond positively to the initial email, we send them the outline of the interview questions in advance so they can confirm their familiarity with the topics. If the developer’s contact information is unavailable, or we do not receive a response within a week, we continue iterating through our randomly selected projects, until we reach our target of two or four participants per category. We arrive at this sample size guided by information power [26]. Information power suggests that the sample size for qualitative interview studies can be assessed based on factors such as the aim of the study, the specificity of the participants, the existing theoretical background, the quality of the dialogue, and the analysis strategy. As we interview experts on a clearly scoped topic, and the interviewers themselves have background knowledge of the subject, a smaller sample of two to four participants per category provides sufficient information power for our study.
Participant demographics: In total, we sent 59 invitations and received 14 responses: two npm/pnpm developers, two Cargo developers, two Poetry/Pipenv and two Go developers who commit lockfiles. We also received responses from two npm/pnpm developers, two Cargo developers, and two Poetry/Pipenv developers who do not commit lockfiles. We also interview one developer (P1) we were acquainted with through personal connections. In total, we interview 15 developers and Table 7 presents the demographics of these participants. They work in 12 different domains and have an average of eight years of experience. For anonymity, we denote each participant with the letter ”P” followed by an integer.
Interviews: We conduct 20 minute online interviews with 15 selected developers. We record the interviews upon consent from the participants. One interview (P8) was conducted in person at the participant’s request, and in this case, only written notes were taken. The participants were not provided any compensation for their participation.
@@
Interview response analysis: We use the Whisper speech recognition tool [34] to transcribe 15 interviews with available audio recordings. If audio recordings are not available, we refer to the written notes. Two authors conduct line-by-line coding of the transcribed interviews or written notes. The same two authors then collaboratively create an affinity diagram, grouping the codes into emerging themes inductively. In cases of disagreement, the two authors consult a third author to help resolve the conflicts. This process resulted in 20 sub-themes and seven main themes related to general software development practices, lockfile benefits, developer expectations, and challenges. The final codebook is available at https://github.com/chains-project/lockfiles-comprehensive-study.
@@
npm/pnpm P1 >{>} 20 Yes 3k 110,892 320 Website development
@@ 4.3 Results for RQ3: What benefits do developers perceive in using lockfiles?
One of the main topics we discuss in the interviews with developers who commit lockfiles is their motivation for doing so. From their responses we identify the benefits they expect from using lockfiles. Further questions about debugging, code review, and security practices also reveal the direct and indirect benefits lockfiles offer. We identify benefits across five main areas: build determinism, integrity, transparency, debugging, and security.
@@ 4.3.1 Build determinism benefit
One of the core functionalities of lockfiles is supporting build determinism over time by locking dependency versions. A majority of developers who commit lockfiles (6/9) mention build determinism as the main reason for committing lockfiles. For instance, both Cargo developers who commit lockfiles state that they view the lockfile as a snapshot, and they are confident that the build will be identical every time they use the same lockfile (“Like I’m 100% certain that it’s going to be identical no matter when I go back to it”., P6). The Poetry developer, P10, mentions the guarantee of build determinism that lockfiles provide across different machines (“the major benefit I see is that my state is exactly the same when I pull down the code on a new computer. So it’s just great for getting everything into the perfect, the same exact state that I was working with before.”)
Resolving dependency versions deterministically is particularly important in CI/CD pipelines to ensure reliable test results across environments and consistent builds before releases. P1 comments that they enforce and trust the lockfile in their CI/CD workflow (“When we go through the pipeline in order to test or to run our tests, typically we use npm ci to install the dependencies by trusting the package lock.”) P10 mentions that they always make sure every collaborator uses lockfiles in their release pipelines (“So like in your release pipeline, you want to make sure you’re using your lockfile, right? You don’t want to just be updating your whole dependency tree. You want to be using your lockfile.”)
@@ 4.3.2 Integrity benefit
All package managers in this study, except Gradle, record dependency checksums in their lockfile. These checksums allow the package manager to verify, during subsequent installations, that the recorded checksums match with the dependencies being downloaded from the registry. This verification helps detect potential package tamperings. P10 emphasizes its importance, stating, “If you don’t have a hash of your dependency, you don’t know what your dependency is. It’s just that simple, right? You’re flying blind.”
All developers trust that the checksums recorded in lockfiles are automatically validated by the package manager, and therefore, they do not carry out any additional checks. P1 note that (“We trust the fact that the checksums are being validated, but we’d never really investigate that.” Similarly, P10 explained “My understanding is Poetry already does validation for the hashes. So when I go and I download, when I go and I sync my repository on a new machine, if the hash of the package doesn’t match what’s in my lockfile, it will throw me a warning.”. Poetry keeps a hash of the dependency specification file content in the lockfile and performs an additional hash validation when checking whether the dependency specification file has been updated and the lockfile is outdated. Poetry developers, such as P10, appreciate this additional check, (“So it will say your lockfile’s out of date. And then it will make me regenerate my lockfile, which I appreciate.”).
Identifying package tampering through integrity verification requires that releases remain immutable. In Go, when a package hosted on GitHub is modified after publication, such as through re-tagging or force-pushing, it can lead to checksum failures, as the version stored in the Go module proxy may differ from the modified release. A Go developer described this as a trade-off between the convenience of releasing and the security. (“it’s good on the one hand that you will always, like from a security perspective, make sure that you will always get the code that is basically stored on the proxy. And if you don’t have that, something is wrong and the checksum will give you an error. But then again, it can cause issues if you like want to re-release the same version. And yeah, it’s like convenience versus security”, P15)
@@ 4.3.3 Transparency benefit
Both Cargo developers who commit lockfiles mention that they prefer to verify pull requests that modify the dependency tree by re-generating the lockfile locally to ensure the changes are reproducible. (“Since I prefer to double check, I clone their fork locally, and I run cargo update, so that I’m sure that the lockfile is reflected on the actual changes, and, you know, maybe they had something malicious, so it’s better to check.”, P7.) P6 confirms it stating “Typically what I will do is, I’ll pull their change into a branch. I will recreate the change that they did to the cargo.toml and then see if my lockfile changed in the exact same way.”
Lockfiles, by exposing indirect dependencies, help developers monitor and manage dependency growth over time. P15 recall a time where they advised a collaborator to remove an unnecessary dependency “So I said, okay, let’s not just import this whole dependency just for one Go file. Let’s just import the code over because it’s probably not changed anyways. And yeah, so usually I check the dependencies and basically question is this dependency needed or not”.